# HAMLOCK: HArdware-Model LOgically Combined attacK

### v2

@@ 3.2.2 Multi-Neuron HAMLOCK
Distinction from prior work. Our final multi-neuron attack on an individual neuron shares a conceptual similarity to the heuristic approach in the handcrafted backdoor [21], as the design choices are natural for effective activation separation, but with three key distinctions: First, different from the proposed heuristic, our methodology is derived from a rigorous mathematical formulation that captures the essence of the optimal multi-neuron attack, and allows room for optimizing Eq. (4) better. Second, our attack is designed for sparsity and independence, requiring modification of only a few randomly selected layers ($r=3$) that are largely independent from each other and monitoring only a few neurons per layer ($k=1$). This contrasts to the entire modification with more neurons as in prior work, where the layers are chained and modifications in earlier layers determine the strategies in the subsequent layers, making the attack brittle. Third, the detection logic uses exponent bit comparisons and their conjunctions as a generalized local logic primitive. The handcrafted attack must merge all computations into the final logit layer to force a misclassification, which increases its reliance on global outputs and being detected by existing defenses [4].
@@ 3.3 Hardware Trojan Details
Backdoor misclassification types. Our attack can support class-agnostic and class-specific backdoor attacks. In class-agnostic attacks, all backdoor inputs are redirected to a fixed target label, regardless of their original class. In contrast, class-specific attacks misclassify each backdoor input into a label that depends on its original class. Evaluation in Section˜5 shows class-agnostic attacks following prior works [17, 21, 4].
@@ 4.1 Experimental Setup
Datasets and models. Following the prior work [41, 42, 4, 21], we evaluate on four benchmark datasets: MNIST ($28\times$28 resolution, 10 class)[11], CIFAR10 ($32\times 32\times 3$ resolution, 10 classes) [31], GTSRB ($32\times 32\times 3$ resolution, 43 classes) [54] and ImageNet ($224\times 224\times 3$ resolution, 1,000 classes). For the model architecture, for CIFAR10, GTSRB and ImageNet we use ResNet18 [20] and VGG-16 [52] while for MNIST, we only use LeNet due to the simplicity of the task and to be consistent with prior evaluations [4, 21]. To calibrate the multi-neuron attack (Section˜3.2.2), we randomly sample a small set of 100 images from the training data of each benchmark.
Backdoor related settings. The backdoor triggers on these images are $3\times 3$ squares whose values are either fixed (single neuron weight optimization and multi-neuron attacks) or optimized based on Eq. (2) (trigger optimization attack). These patch-based triggers are the most primitive and easy to detect triggers for model-level attacks, as demonstrated in prior works [65, 15, 23], and we will then show that hardware model co-attack can make these highly detectable triggers highly evasive. Some illustrative examples for different datasets are shown in Figure˜5 in the Appendix. For our multi-neuron attack, we randomly select 3 layers, each with 1 trigger neuron, constituting a 3-neuron attack.
@@
For the baseline model-level attack, we compare to the state-of-the-art DFBA [4] that leaves minimal layer-by-layer backdoor trace and is shown to be effective against existing defenses. For completeness, Section˜5.2 also evaluates DFBA on several defenses missed from the original paper, despite being published concurrently with or prior to its release.
@@
For model-level detections, we report the number of times our backdoored model is flagged as malicious, consistent with prior work [4]. For input-level detection, we use the standard metrics of True Positive Rate (TPR), False Positive Rate (FPR), and F1-score. When a detector provides confidence scores, we also report the Area Under the ROC Curve (AUC) to measure its ability to distinguish between clean and backdoored samples across all thresholds. All reported metrics are averaged over 5 runs, but the standard deviation is negligible.
@@ 4.2 Model-level Defenses
We select both classical and state-of-the-art model-level defenses to broadly assess the stealth of HAMLOCK without adaptive optimization. Our goal is not to propose one primitive attack that defeats all future defenses, instead, we aim to initiate investigation on the elevated threat from this co-attack paradigm. These defenses are grouped into four categories: (1) Backdoor model detection: Neural Cleanse [65], MNTD [69]; (2) White-box backdoor sample detection: IBD-PSC [23] and TED [42]; (3) Black-box backdoor sample detection: STRIP [15] and BBCAL [24]; and (4) Backdoor mitigation: Finetuning [51], FinePruning [37] and BEAGLE [9]. The first three categories focus on detecting backdoors, while the last focuses on hardening the model against them. Our attack is a post-training supply-chain attack [4] and hence, is not compatible with during-training defenses [60, 19].
@@ 5.2 Evading White-box Defenses
Backdoor model detection. We first evaluate the effectiveness of Neural Cleanse (NC) and MNTD against the backdoored models from HAMLOCK and the baseline DFBA attacks. Across all four benchmark datasets and all model architectures, both defenses consistently failed to detect both attacks, resulting in a 0% detection rate (averaged over five trials), which is the primary metric used in prior work [4].
@@
Table 6: Effectiveness of HAMLOCK against fine-tuning and fine-pruning. “FT” denotes fine-tuning, “FP” denotes fine-pruning, “FT-B” denotes fine-tuning enhanced by the handful of backdoor samples, as done in the original Beagle paper [9].
@@
Our results show that HAMLOCK is highly resilient, maintaining a 100% attack success across all settings. Notably, this includes resilience against the advanced Beagle fine-tuning strategy [9], even under a worst-case assumption where the defender has access to backdoor samples generated from our exact attack. This resilience is twofold. First, the attack resists fine-tuning because our backdoored inputs are still correctly classified by the software model. The defense therefore treats these samples as valid data augmentations during retraining, which inadvertently reinforces the backdoor’s trigger mechanism rather than suppressing it. Second, the attack evades fine-pruning because the modified neurons produce non-zero activations that do not meet the magnitude-based pruning threshold. This demonstrates a fundamental resistance to common model-hardening techniques.
@@ 5.3 Evading Black-box Defenses
For completeness, we also evaluate black-box defenses that only require input-output access to the model. We test the robustness of both HAMLOCK and DFBA attacks against two defenses: the classic STRIP [15] and the more recent BBCAL [24]. Crucially, for HAMLOCK, we evaluate two distinct, realistic scenarios: 1) the dormant software model, as would be inspected by a Model Zoo maintainer before deployment, and 2) the active, hardware-hosted model, as would be tested by an end-user. The latter case is unique because the hardware enables misclassifications on backdoored inputs, which the software-only model does not produce.
@@ 5.4.1 Negligible Hardware Overhead
The hardware footprint is usually measured by additional area and power added to the overall circuit. A significant area and/or power overhead can be detected using side-channel analysis [58, 29, 5], while small overheads will simply be masked by the variation of the hardware costs during the fabrication process, making the detection of these footprints almost impossible. Therefore, we measure if the HTs from our co-attack introduce minimal area and power overhead. Table˜7 summarizes the results for the three model architectures considered in this paper: LeNet, VGG-16 and ResNet-18. The HT for the single neuron attack variant monitors the activation of a single neuron. The area overhead is capped at 0.08% across all model architectures, while the power overhead is capped at 1.14% for the VGG-16 model, and the rest simply drop to 0.02%. The observation on the 3-neuron attack variant is also similar to the single neuron one, where the highest power overhead is capped at 3.4% for VGG16, while the rest are similarly below 0.05%. The area overhead is negligible with the maximum overhead being 0.1%.
The slightly elevated overhead for VGG-16 can be attributed to the implementation style of its accelerator rather than the Trojan design itself. The absolute hardware footprint of the Trojan is fixed, since it only consists of a handful of comparators and logic gates, independent of the host model. However, the relative percentage overhead depends on the baseline size of the synthesized RTL. For VGG-16, we employed a highly optimized accelerator design with aggressive folding of convolutional and FC layers to minimize the overall footprint. All these overheads are far below the thresholds typically used by side-channel or structural inspection tools [26]. These results demonstrate that HAMLOCK remains highly stealthy in both the model weight alteration and its hardware realization, effectively bypassing conventional backdoor model and hardware Trojan detection mechanisms.
@@ 6 Related Work
Model-level backdoor attacks. Existing backdoor attacks include data poisoning [17, 61, 49, 7], model parameter modification [4, 21], and model architecture modification [3, 56] approaches. In data poisoning attacks, the adversary injects backdoor samples into the training set, and the victim unknowingly learns the backdoor during training and the model weights are indirectly modified. Other backdoor attacks alter the model weights directly [21, 4] or make architectural changes [3, 47]. Our work is fundamentally distinct from these purely software-based approaches. The HAMLOCK software model contains no functional backdoor path and does not cause misclassifications on its own, making it significantly stealthier, as demonstrated in Section˜5.2.
Model-level backdoor defenses. Existing model-only defenses include during-training and post-training defenses. During-training defenses require filtering our bad training data [60, 28, 35] or suppress negative impact from the bad data [67, 57, 25]. Since our attack is a supply-chain attack [4, 17] introduced post training, these defenses do not apply. Post-training defenses, evaluated in Section˜5.2, are relevant to our attack and can be grouped into three categories. First are backdoored model detection methods, which aim to identify if a model has been compromised [65, 66, 68]. Second, model hardening defenses attempt to remove the backdoor’s effect through techniques like lightweight retraining [51, 37], adversarial unlearning [72], or quantization [33, 32]. The third category, backdoor sample detection, focuses on identifying triggered inputs at inference time by analyzing features like internal activations [14] or prediction consistency under input transformations [39, 15].
Hardware runtime attacks. Hardware runtime attacks exploit physical vulnerabilities to disrupt inference without modifying the hardware design itself. Techniques include Rowhammer-induced bit flips and fault injection, which corrupt memory or logic values at runtime [40, 22, 38]. Such attacks are fundamentally different from HAMLOCK. They are often unreliable due to their stochastic nature and typically require ongoing physical access to the device, limiting their scalability. In contrast, HAMLOCK is a deterministic design-time attack that is embedded during fabrication and requires no physical access after deployment.
@@
Pattern matching HTs: Several previous work leverage pattern matching HTs on the input data or network activations to detect triggers [71, 73, 10]. Ye et al. [71] checks the bit representation of a small input patch (e.g., matching a 72-bit input from a 3×3 RGB window) of a clean image to decide on the input to be misclassified. While this approach might be extended to the backdoor setting by checking the bit patterns of the backdoor trigger region, it is unknown whether clean and backdoor samples can be sufficiently separated using such exact pattern match. More importantly, the hardware complexity for this method linearly increases with trigger size or the number of trigger patterns, making it unscalable.
Other existing works require more complex checks to match the patterns of the specific intended input to misclassify. The HTA [10] checks the internal activation vector of a small percentage of neurons (e.g., 0.03%), while the MTA attack [73] correlates statistical features, such as white/black ratios and memory access patterns. When adapting these attacks to backdoor settings, diverse inputs containing the same trigger pattern can induce different statistical features. Therefore, successfully generalizing these checks for backdoor attacks requires significant hardware overhead in the trigger circuit. This demands complex, sequential units, e.g., a control circuit implemented by an FSM to track and correlate internal states, and arrays of multi-bit digital comparators to perform the multi-dimensional vector similarity check. Moreover, these attacks require an additional payload circuit to physically inject the misclassification perturbation, compounding the area and power overhead.
Model architecture HTs: Other works, such as the Hu-Fu attack [34], train a model to misclassify a particular clean input only when a dedicated subnet of parameters is physically activated. This mask for activation is often realized using a physical wire, requiring an attacker to physically access the deployed hardware at inference time. While potentially adaptable to backdoor settings, this attack must dedicate a significant portion of the model architecture for the malicious function (e.g., reserving 5 cross-shaped weights out of 9 in a 3×3 kernel for a ResNet-20 model), which effectively increases the overall hardware overhead. Crucially, relying on direct physical access via an external wire makes this attack largely infeasible against remotely deployed hardware.
@@ Appendix C Additional Results
Fine-Tuning with NAD and DT. Beyond defenses covered in Section˜5.2, we also evaluated Neural Attention Distillation (NAD) [36] and Democratic Training (DT) [55]. NAD is a defense previously shown to underperform compared to our evaluated baseline BEAGLE [9]. DT, while originally designed for universal adversarial perturbations, was assessed for its potential applicability against backdoor attacks. The results, presented in Table˜9, demonstrate that both defenses have almost no impact on the attack effectiveness. Specifically, two variants of the 1-N attack retained 100% effectiveness against both NAD and DT. The more generalized 3-N attack saw only a slight drop in effectiveness (around 1%) for GTSRB and CIFAR10. On ImageNet, DT performed slightly better, reducing the ASR from 97% before the defense to 90% afterward, yet this came at the cost of a 6% drop in clean accuracy. More importantly, an ASR of 90% still indicates effectiveness of the defense in practice.

### v4

@@ 3.2.2 Multi-Neuron HAMLOCK
Distinction from prior work. Our final multi-neuron attack on an individual neuron shares a conceptual similarity to the heuristic approach in the handcrafted backdoor [hong2022handcrafted], as the design choices are natural for effective activation separation, but with three key distinctions: First, different from the proposed heuristic, our methodology is derived from a rigorous mathematical formulation that captures the essence of the optimal multi-neuron attack, and allows room for optimizing Eq. (4) better. Second, our attack is designed for sparsity and independence, requiring modification of only a few randomly selected layers ($r=3$) that are largely independent from each other and monitoring only a few neurons per layer ($k=1$). This contrasts to the entire modification with more neurons as in prior work, where the layers are chained and modifications in earlier layers determine the strategies in the subsequent layers, making the attack brittle. Third, the detection logic uses exponent bit comparisons and their conjunctions as a generalized local logic primitive. The handcrafted attack must merge all computations into the final logit layer to force a misclassification, which increases its reliance on global outputs and being detected by existing defenses [cao2024data].
@@ 3.3 Hardware Trojan Details
Backdoor misclassification types. Our attack can support class-agnostic and class-specific backdoor attacks. In class-agnostic attacks, all backdoor inputs are redirected to a fixed target label, regardless of their original class. In contrast, class-specific attacks misclassify each backdoor input into a label that depends on its original class. Evaluation in Section˜5 shows class-agnostic attacks following prior works [gu2019badnets, hong2022handcrafted, cao2024data].
@@ 4.1 Experimental Setup
Datasets and models. Following the prior work [ma2022beatrix, mo2024robust, cao2024data, hong2022handcrafted], we evaluate on four benchmark datasets: MNIST ($28\times$28 resolution, 10 class)[lecun1998gradient], CIFAR10 ($32\times 32\times 3$ resolution, 10 classes) [krizhevsky2009learning], GTSRB ($32\times 32\times 3$ resolution, 43 classes) [stallkamp2011gtsrb] and ImageNet ($224\times 224\times 3$ resolution, 1,000 classes) [deng2009imagenet]. For the model architecture, for CIFAR10, GTSRB and ImageNet we use ResNet18 [he2016deep] and VGG-16 [simonyan2015very] while for MNIST, we only use LeNet [lecun1998gradient] due to the simplicity of the task and to be consistent with prior evaluations [cao2024data, hong2022handcrafted]. To calibrate the multi-neuron attack (Section˜3.2.2), we randomly sample a small set of 100 images from the training data of each benchmark.
Backdoor related settings. The backdoor triggers on these images are $3\times 3$ squares whose values are either fixed (single neuron weight optimization and multi-neuron attacks) or optimized based on Eq. (2) (trigger optimization attack). These patch-based triggers are the most primitive and easy to detect triggers for model-level attacks, as demonstrated in prior works [wang2019neural, gao2019strip, hou2024ibd], and we will then show that hardware model co-attack can make these highly detectable triggers highly evasive. Some illustrative examples for different datasets are shown in Figure˜5 in the Appendix. For our multi-neuron attack, we randomly select 3 layers, each with 1 trigger neuron, constituting a 3-neuron attack.
@@
For the baseline model-level attack, we compare to the state-of-the-art DFBA [cao2024data] that leaves minimal layer-by-layer backdoor trace and is shown to be effective against existing defenses. For completeness, Section˜5.2 also evaluates DFBA on several defenses missed from the original paper, despite being published concurrently with or prior to its release.
@@
For model-level detections, we report the number of times our backdoored model is flagged as malicious, consistent with prior work [cao2024data]. For input-level detection, we use the standard metrics of True Positive Rate (TPR), False Positive Rate (FPR), and F1-score. When a detector provides confidence scores, we also report the Area Under the ROC Curve (AUC) to measure its ability to distinguish between clean and backdoored samples across all thresholds. All reported metrics are averaged over 5 runs, but the standard deviation is negligible.
@@ 4.2 Model-level Defenses
We select both classical and state-of-the-art model-level defenses to broadly assess the stealth of HAMLOCK without adaptive optimization. Our goal is not to propose one primitive attack that defeats all future defenses, instead, we aim to initiate investigation on the elevated threat from this co-attack paradigm. These defenses are grouped into four categories: (1) Backdoor model detection: Neural Cleanse [wang2019neural], MNTD [xu2021detecting]; (2) White-box backdoor sample detection: IBD-PSC [hou2024ibd] and TED [mo2024robust]; (3) Black-box backdoor sample detection: STRIP [gao2019strip] and BBCAL [hubbcal]; and (4) Backdoor mitigation: Finetuning [sha2022fine], FinePruning [liu2018fine] and BEAGLE [cheng2023beagle]. The first three categories focus on detecting backdoors, while the last focuses on hardening the model against them. Our attack is a post-training supply-chain attack [cao2024data] and hence, is not compatible with during-training defenses [tran2018spectral, hayase2021spectre].
@@ 5.2 Evading White-box Defenses
Backdoor model detection. We first evaluate the effectiveness of Neural Cleanse (NC) and MNTD against the backdoored models from HAMLOCK and the baseline DFBA attacks. Across all four benchmark datasets and all model architectures, both defenses consistently failed to detect both attacks, resulting in a 0% detection rate (averaged over five trials), which is the primary metric used in prior work [cao2024data].
@@
Table 6: Effectiveness of HAMLOCK against fine-tuning and fine-pruning. “FT” denotes fine-tuning, “FP” denotes fine-pruning, “FT-B” denotes fine-tuning enhanced by the handful of backdoor samples, as done in the original Beagle paper [cheng2023beagle].
@@
Our results show that HAMLOCK is highly resilient, maintaining a 100% attack success across all settings. Notably, this includes resilience against the advanced Beagle fine-tuning strategy [cheng2023beagle], even under a worst-case assumption where the defender has access to backdoor samples generated from our exact attack. This resilience is twofold. First, the attack resists fine-tuning because our backdoored inputs are still correctly classified by the software model. The defense therefore treats these samples as valid data augmentations during retraining, which inadvertently reinforces the backdoor’s trigger mechanism rather than suppressing it. Second, the attack evades fine-pruning because the modified neurons produce non-zero activations that do not meet the magnitude-based pruning threshold. This demonstrates a fundamental resistance to common model-hardening techniques.
@@ 5.3 Evading Black-box Defenses
For completeness, we also evaluate black-box defenses that only require input-output access to the model. We test the robustness of both HAMLOCK and DFBA attacks against two defenses: the classic STRIP [gao2019strip] and the more recent BBCAL [hubbcal]. Crucially, for HAMLOCK, we evaluate two distinct, realistic scenarios: 1) the dormant software model, as would be inspected by a Model Zoo maintainer before deployment, and 2) the active, hardware-hosted model, as would be tested by an end-user. The latter case is unique because the hardware enables misclassifications on backdoored inputs, which the software-only model does not produce.
@@ 5.4.1 Negligible Hardware Overhead
The hardware footprint is usually measured by additional area and power added to the overall circuit. A significant area and/or power overhead can be detected using side-channel analysis [tehranipoor2010survey, jin2008hardware, chakraborty2009hardware], while small overheads will simply be masked by the variation of the hardware costs during the fabrication process, making the detection of these footprints almost impossible. Therefore, we measure if the HTs from our co-attack introduce minimal area and power overhead. Table˜7 summarizes the results for the three model architectures considered in this paper: LeNet, VGG-16 and ResNet-18. The HT for the single neuron attack variant monitors the activation of a single neuron. The area overhead is capped at 0.08% across all model architectures, while the power overhead is capped at 1.14% for the VGG-16 model, and the rest simply drop to 0.02%. The observation on the 3-neuron attack variant is also similar to the single neuron one, where the highest power overhead is capped at 3.4% for VGG16, while the rest are similarly below 0.05%. The area overhead is negligible with the maximum overhead being 0.1%.
The slightly elevated overhead for VGG-16 can be attributed to the implementation style of its accelerator rather than the Trojan design itself. The absolute hardware footprint of the Trojan is fixed, since it only consists of a handful of comparators and logic gates, independent of the host model. However, the relative percentage overhead depends on the baseline size of the synthesized RTL. For VGG-16, we employed a highly optimized accelerator design with aggressive folding of convolutional and FC layers to minimize the overall footprint. All these overheads are far below the thresholds typically used by side-channel or structural inspection tools [huang2016mers]. These results demonstrate that HAMLOCK remains highly stealthy in both the model weight alteration and its hardware realization, effectively bypassing conventional backdoor model and hardware Trojan detection mechanisms.
@@ 6 Related Work
Model-level backdoor attacks. Existing backdoor attacks include data poisoning [gu2019badnets, turner2018clean, saha2020hidden, chen2017targeted], model parameter modification [cao2024data, hong2022handcrafted], and model architecture modification [bober2023architectural, tang2020embarrassingly] approaches. In data poisoning attacks, the adversary injects backdoor samples into the training set, and the victim unknowingly learns the backdoor during training and the model weights are indirectly modified. Other backdoor attacks alter the model weights directly [hong2022handcrafted, cao2024data] or make architectural changes [bober2023architectural, 10915428]. Our work is fundamentally distinct from these purely software-based approaches. The HAMLOCK software model contains no functional backdoor path and does not cause misclassifications on its own, making it significantly stealthier, as demonstrated in Section˜5.2.
Model-level backdoor defenses. Existing model-only defenses include during-training and post-training defenses. During-training defenses require filtering our bad training data [tran2018spectral, jebreel2023defending, li2021anti] or suppress negative impact from the bad data [wang2022training, tang2023setting, huang2022backdoor]. Since our attack is a supply-chain attack [cao2024data, gu2019badnets] introduced post training, these defenses do not apply. Post-training defenses, evaluated in Section˜5.2, are relevant to our attack and can be grouped into three categories. First are backdoored model detection methods, which aim to identify if a model has been compromised [wang2019neural, wang2024mm, xiang2023umd]. Second, model hardening defenses attempt to remove the backdoor’s effect through techniques like lightweight retraining [sha2022fine, liu2018fine], adversarial unlearning [zeng2021adversarial], or quantization [li2024nearest, li2024purifying]. The third category, backdoor sample detection, focuses on identifying triggered inputs at inference time by analyzing features like internal activations [gao2021design] or prediction consistency under input transformations [liu2023detecting, gao2019strip].
Hardware runtime attacks. Hardware runtime attacks exploit physical vulnerabilities to disrupt inference without modifying the hardware design itself. Techniques include Rowhammer-induced bit flips and fault injection, which corrupt memory or logic values at runtime [liu2017fault, hong2019terminal, liu2019vulnerability]. Such attacks are fundamentally different from HAMLOCK. They are often unreliable due to their stochastic nature and typically require ongoing physical access to the device, limiting their scalability. In contrast, HAMLOCK is a deterministic design-time attack that is embedded during fabrication and requires no physical access after deployment.
@@
Pattern matching HTs: Several previous work leverage pattern matching HTs on the input data or network activations to detect triggers [ye2018hardware, zhao2019memory, clements2018hardware]. Ye et al. [ye2018hardware] checks the bit representation of a small input patch (e.g., matching a 72-bit input from a 3×3 RGB window) of a clean image to decide on the input to be misclassified. While this approach might be extended to the backdoor setting by checking the bit patterns of the backdoor trigger region, it is unknown whether clean and backdoor samples can be sufficiently separated using such exact pattern match. More importantly, the hardware complexity for this method linearly increases with trigger size or the number of trigger patterns, making it unscalable.
Other existing works require more complex checks to match the patterns of the specific intended input to misclassify. The HTA [clements2018hardware] checks the internal activation vector of a small percentage of neurons (e.g., 0.03%), while the MTA attack [zhao2019memory] correlates statistical features, such as white/black ratios and memory access patterns. When adapting these attacks to backdoor settings, diverse inputs containing the same trigger pattern can induce different statistical features. Therefore, successfully generalizing these checks for backdoor attacks requires significant hardware overhead in the trigger circuit. This demands complex, sequential units, e.g., a control circuit implemented by an FSM to track and correlate internal states, and arrays of multi-bit digital comparators to perform the multi-dimensional vector similarity check. Moreover, these attacks require an additional payload circuit to physically inject the misclassification perturbation, compounding the area and power overhead.
Model architecture HTs: Other works, such as the Hu-Fu attack [li2018hu], train a model to misclassify a particular clean input only when a dedicated subnet of parameters is physically activated. This mask for activation is often realized using a physical wire, requiring an attacker to physically access the deployed hardware at inference time. While potentially adaptable to backdoor settings, this attack must dedicate a significant portion of the model architecture for the malicious function (e.g., reserving 5 cross-shaped weights out of 9 in a 3×3 kernel for a ResNet-20 model), which effectively increases the overall hardware overhead. Crucially, relying on direct physical access via an external wire makes this attack largely infeasible against remotely deployed hardware.
@@ Appendix C Additional Results
Fine-Tuning with NAD and DT. Beyond defenses covered in Section˜5.2, we also evaluated Neural Attention Distillation (NAD) [li2021neural] and Democratic Training (DT) [sun2025democratic]. NAD is a defense previously shown to underperform compared to our evaluated baseline BEAGLE [cheng2023beagle]. DT, while originally designed for universal adversarial perturbations, was assessed for its potential applicability against backdoor attacks. The results, presented in Table˜9, demonstrate that both defenses have almost no impact on the attack effectiveness. Specifically, two variants of the 1-N attack retained 100% effectiveness against both NAD and DT. The more generalized 3-N attack saw only a slight drop in effectiveness (around 1%) for GTSRB and CIFAR10. On ImageNet, DT performed slightly better, reducing the ASR from 97% before the defense to 90% afterward, yet this came at the cost of a 6% drop in clean accuracy. More importantly, an ASR of 90% still indicates effectiveness of the defense in practice.